OpenStudy (anonymous):

which of the following could be the graph of the y= -3x - 2

OpenStudy (denonakavro):



OpenStudy (shouborno):

 \(y=-3x-2\) \(or, \frac x {-\frac {2}3} +\frac y{-2} = 1 \) From this format, we can derive the intercepting points on the two axis, and thus draw the line easily.